# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

### v1

@@ 3. An Overview of AutoPrep
Figure 4. Our proposed Multi-agent Data Prep Framework for TQA (Left), which could be plugged to any other TQA method (Right), such as NL2SQL as shown in the Figure. The proposed framework AutoPrep decomposes the data prep process into two stages: ① The Planning stage and ② the Executing stage.
@@
    Multiple Programmer Agents (i.e., [Uncaptioned image] Augmenter): the Programming stage. It directs the LLM to generate low-level implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$.
@@
    An [Uncaptioned image] Executor Agent: the Executing stage. It executes the given low-level code and debugs iteratively to improve the generated code quality.
@@
  • $T^{\prime}=$Filter($T$, [Date, Cyclist, Medal]) that filters out column Age, which is irrelevant to the question;
@@
(c) The Executing stage: a Executor agent iteratively refines the generated code if any errors occur.
@@
The Planner agent and the three Programmer agents (i.e., Augmenter, Normalizer, and Filter) are the core modules of AutoPrep, which will be discussed in more details below.
The Planner Agent. Given a question $Q$ and a table, the Planner agent generates a sequence of high-level operations for the specific question, as illustrated in the previous example. The key challenge here is how to suggest question-aware operations that address specific NL questions. Specifically, even for the same table, different NL questions may require not only different data prep operations but also varying sequences of those operations. For example, the second Normalize operation might not be necessary if “in February” is not part of the question. To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent. This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch. Compared to existing CoT methods (Wei et al., 2022), which break down questions into sub-questions, our approach better captures the relationships between questions and tables. More details of the CoC method are given in Section 4.
The Programmer Agents. Given a high-level operation $o_{i}$, a Programmer agent is responsible for generating low-level code $f_{i}$ to execute the operation. As discussed previous, we develop specialized agents i.e., Augmenter, Normalizer, and Filter for various types of operations. The key challenge is that a given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial. For example, the Augmenter agent may generate an overly generic regular expression that extracts countries based on parentheses. Unfortunately, this code fails to correctly process “Dav. ITA”, which is formatted differently from other values. To tackle this challenge, we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions. We establish a search space of API functions for all high-level operations, create an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism. More details of our tool-augmented approach are discussed in Section 5.
@@
4. The Planner Agent
Given a question $Q$ posed over a table $T$, Planner aims to generate a logical plan that outlines a sequence of high-level operations needed to prepare the table for answering the question. In this section, we first introduce a direct prompting method in Section 4.1 that better demonstrates the key challenge in designing the Planner agent. Then, to address the challenge, we propose our Chain-of-Clauses (CoC) reasoning method in Section 4.2.
@@ 4.1. A Direct Prompting Method
The most common way to generate a logical plan is to directly prompt an LLM using a typical in-context learning approach, as shown in Algorithm 1. The inputs are a question $Q$, a table $T$, a set $\Sigma$ of specifications for each operation type, and an LLM $\theta$. Here, each specification $\sigma\in\Sigma$ describes the purpose of an operation type, e.g., “an Augment operation augments a new column for a table based on existing columns, in response to the specific needs of a question”. The output of the algorithm is a set $O$ of high-level operations like those shown in Figure 4, e.g., Normalize(“Cast to INT”, Medal).
Algorithm 1 Direct Prompting for the Planner Agent
1:A question $Q$, a table $T$, an operation specification set $\Sigma$, and a large language model (LLM) $\theta$.
2:// Prepare $k$ in-context exemplars (i.e., demos)
3:$\mathcal{D}=\{(T_{i},Q_{i},\Sigma,O_{i})\}_{i=1}^{k}$
4:// Prompt the LLM for inferring the operations
5:$O\leftarrow\theta(T,Q,\Sigma,\mathcal{D})$
6:return $O$
Specifically, Algorithm 1 first prepares a set $\mathcal{D}$ of $k$ exemplars $\{(T_{i},Q_{i},\Sigma,O_{i})\}_{i=1}^{k}$ that demonstrates to LLM $\theta$ the goal of this logical plan generation task. Then, the algorithm prompts the LLM to infer the operations $O$ for our input table $T$ and question $Q$.
@@ 4.2. A Chain-of-Clauses Method
To address this challenge, we propose a Chain-of-Clauses (CoC) reasoning method the decomposes the entire process of logical plan generation into two phrases, as shown in Figure 5(b).
@@
Compared with existing CoT methods (Wei et al., 2022), which simply break down questions into sub-questions, our approach is more effective for logical plan generation due to the following reasons. First, our method decomposes the question into a set of analysis steps over the table, formalized as an SQL-like Analysis Sketch, which simplifies the task of logical plan generation. More importantly, our method jointly considers each analysis step along with the corresponding relevant data to prompt the LLM, effectively capturing the relationships between the question and the table.
Algorithm 2 CoC Reasoning for the Planner Agent
1:A question $Q$, a table $T$, an operation specification set $\Sigma$, and a large language model (LLM) $\theta$.
2:$O\leftarrow\emptyset$
3:// Phase I: Analysis Sketch generation
4:$S\leftarrow\theta(Q,T,\{(Q_{i},T_{i},s_{i})\})$
5:// Phase II: High-level operation generation
6:for each clause $s$ in $S$ do
7:    $T^{s}\leftarrow$ extract_data($T$)
8:    $o\leftarrow\theta(s,T^{s},\Sigma,\{s_{i},T^{s}_{i},\Sigma,o_{i}\})$
9:    $O\leftarrow O\cup\{o\}$
10:end for
11:return $O$
Specifically, Algorithm 2 presents the pseudo-code of our CoC reasoning method, which takes as input a question $Q$, a table $T$, a specification set $\Sigma$, and an LLM $\theta$ to produce an operation set $O$. To this end, we introduce a key concept of SQL-like Analysis Sketch, which is formally defined as follows.
@@
where ${\tt agg}(A)$ is an aggregation function (e.g., SUM and AVG) over column $A$, ${\tt Pred}(A_{i})$ denotes a predicate (i.e., filtering condition) over column $A_{i}$, such as Date LIKE02-%. Note that $f(\{A_{i}\})$ is a user-defined function (UDF) that maps existing columns $\{A_{i}\}$ in the table into a new column. Figure 5(b) shows an example Analysis Sketch with a UDF function that specifies a new column Country, i.e., $f(T)$ AS Country. Obviously, this UDF is introduced because the Analysis Sketch contains a GROUP BY Country clause.
Given the key concept Analysis Sketch, Algorithm 2 produces a logical plan via two phases, as shown in the following example.
@@
  • $f(T)$ AS Country: this clause and the entire table $T$ are used to prompt the LLM to generate an Augment operation.
@@
  • Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize operation.
Finally, the algorithm prompt the LLM to generate a Filter operation that only selects the columns relevant to the Analysis Sketch.
5. The Programmer & Executor Agents
@@
The Programmer agents translate a high-level logical plan into a physical plan by generating low-level code, which is then passed to an Executor agent for iterative code execution and debugging. A straightforward approach is to prompt an LLM with the logical plan using typical in-context learning with several exemplars, asking the LLM to generate code for each high-level operation in the plan. The code is then executed iteratively, and if any runtime errors occur, the Debugger is prompted with the error messages for debugging.
However, the above method may have limitations, as the generated code may be overly generic and unable to effectively address the heterogeneity challenge of the tables. Specifically, many tables originate from web sources or real-world scenarios, where values have diverse syntactic (e.g., “19-Oct” and “9/14” in $T_{6}$) or semantic formats (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code customized to these tables. For example, consider table $T_{4}$ in 2(b). Given a Normalize operation to standardize country formats, a generic function from an existing Python library may not be sufficient to transform country names (e.g., “Italia”) to their ISO codes (e.g., “ITA”). In such cases, customized functions, such as the clean_country function from an external library dataprep (Peng et al., 2021), are needed to address specific requirements of code generation.
To address these limitations, we introduce a tool-augmented method for the Programmer agents, enhancing the LLM’s code generation capabilities by utilizing pre-defined API functions, referred to as the function pool in this paper. Specifically, we first introduce the key idea of our tool-augmented method in  5.1, then present the function pool in Section 5.2, and finally develop an algorithm for code generation and execution in Section 5.3.
5.1. Tool-Augmented Method: Key Idea
Figure 6 provides an overview of our tool-augmented method for physical plan generation and execution. Given a table $T$ and a set $O={o_{1},\ldots,o_{|O|}}$ of high-level operations, the method iteratively generates and executes low-level code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$.
@@ 5.1. Tool-Augmented Method: Key Idea
Code Execution and Debugging. Given the function generated in the previous step, the Executor agent then applies the function over table $T_{i}$ and support iterative debugging if there are any runtime errors or incorrect results.
For instance, as shown in Figure 6, the first lambda function for the argument func only extracts countries based on parentheses, which may produce incorrect results for the value “Dav.ITA”, as it is formatted differently from other values. In this case, based on the execution results, the Executor agent will refine the regular expression in the lambda function until no runtime errors or incorrect results occur.
After completing these two steps, the method produces a new intermediate table $T_{i+1}$ and proceeds to the next operation, $o_{i+1}$, i.e., the Normalize operation, as shown in Figure 6.
@@ 5.2. Search Space of Function Pools
This section presents a search of function pools for the different types of high-level operations formalized in the current AutoPrep.
@@
  • $\mathtt{extract}$. This function extracts a substring from a column to generate a new column. It has two arguments: ${\tt column}$, representing the name of the source column, and ${\tt func}$, representing the substring extraction function. Using table $T_{1}$ from Figure 2(a) as an example, to extract the country code, we would set ${\tt column}$ to Cyclist and ${\tt func}$ to the lambda function lambda x: re.search(r’(.?)(.*?)( . ∗ ? )’,x).group(1).
  • $\mathtt{calculate}$. This function generates a new column through arithmetic operations of columns such as addition, subtraction, multiplication, or a combination of these. To use this operator, we need to provide two arguments: ${\tt columns}$ and ${\tt func}$. The argument ${\tt columns}$ should be a list of source column names. Similar to $\mathtt{extract}$, the ${\tt func}$ argument is typically a lambda function. However, for $\mathtt{calculate}$, the input to the function should be a dictionary of values rather than a string, since we may perform calculations on multiple columns. For example, given a table with two columns A and B, representing the scores of team A and team B, we need to generate a new column ScoreDifference. In this case, we set ${\tt columns}$ to [’A’, ’B’] and ${\tt func}$ to the specific lambda function lambda x: x[’A’]-x[’B’].
  • $\mathtt{map\mbox{-}to\mbox{-}boolean}$ and $\mathtt{concatenate}$. Both functions generate a new column from multiple columns. The difference is that one generates boolean values, while the other concatenates new strings. Both operators take two arguments, ${\tt columns}$ and ${\tt func}$, similar to those in the $\mathtt{calculate}$ operator.
@@
  • $\mathtt{to\mbox{-}numerical}$. This function standardizes a column to a numerical type, such as int or float. To use it, two arguments need to be specified: ${\tt column}$, representing the source column, and ${\tt func}$, representing a lambda function. For example, given the table $T_{5}$ in Figure 2(b), we use this function to standardize the Females column to integers, with ${\tt column}$ set to Females and ${\tt func}$ as lambda x: int(x.replace(’,’, ’’)).
@@
  • $\mathtt{clean\mbox{-}string}$. To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$. It has two arguments: ${\tt column}$ and ${\tt trans_{d}ict}$. The $\tt trans_{d}ict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$. Given the original table $T_{4}$ in Figure  2(b), we set $trans_{d}ict$ to {’Italia’: ’ITA’}.
@@
Discussions. It is important to note that the search space is extensible, allowing additional functions or external APIs to be easily incorporated into the function pools. Moreover, AutoPrep can also include external API functions, such as specific data prep functions for various semantic types (e.g., date, name, company) developed in the dataprep library (Peng et al., 2021).
5.3. Code Generation and Execution
Algorithm 3 Low-level Code Generation and Execution
1:A question $Q$, a table $T$, a number $r$, a high-level operation set $O$ and three function pools $\mathcal{F}_{\tt Aug}$, $\mathcal{F}_{\tt Norm}$ and $\mathcal{F}_{\tt Filter}$, an LLM $\theta$
2:$F\leftarrow\varnothing,~{}T_{1}\leftarrow T$
3:for each operation $o_{i}$ in $O$ do
4:    // Step 1: Function selection and argument inference
5:    $\mathcal{F}\leftarrow$ Assign($o_{i}$, [$\mathcal{F}_{\tt Aug}$, $\mathcal{F}_{\tt Norm}$, $\mathcal{F}_{\tt Filter}$])
6:    $f_{i}\leftarrow\theta(o_{i},Q,T,\mathcal{F})$
7:    // Step 2: Code generation and execution
8:    while $f_{i}$ has errors &\&& ${\tt debug\_round}<r$ do
9:         $f_{i}\leftarrow$ Debug($Q,T,f_{i}$)
[…]
11:end for
12:return $T_{|O|+1}$
Algorithm 3 presents the pseudo-code of our tool-augmented approach for code generation and execution. The algorithm takes as input a question $Q$, a table $T$, a set $O$ of high-level operations, three function pools, and an LLM, and produces a prepared table $T_{|O|+1}$. The algorithm iteratively examines the operations in $O$. In each iteration, it first generates a function from the corresponding function pool and refines it through multiple rounds of debugging. The algorithm then applies the generated function to the corresponding intermediate table and proceeds to the next iterations.
@@ 6.1. Experimental Setup
Datasets. We evaluate our proposed framework AutoPrep using two well-adopted benchmarking datasets WikiTQ (Pasupat and Liang, 2015) and TabFact (Chen et al., 2019). The statistics are shown in Table 1.
  • WikiTQ. WikiTableQuestions (WikiTQ) is a dataset designed for QA tasks, containing complex questions annotated by crowd workers based on diverse Wikipedia tables. WikiTQ comprises 17,689 question-answer pairs in the training set and 4,344 in the test set with significant variation in table size. The answers in WikiTQ can be categorized into three forms: (i) string from a cell in the table, (ii) string from analyzing, and (iii) list of cells in the table.
  • TabFact. Table-Fact-Checking (TabFact) dataset provides a collection of Wikipedia tables, along with manually annotated NL statements. One requires to deduce the relations between statements and tables as “true” (a statement is entailed by a table) or “false” (a statement if refuted by a table). To reduce the experimental cost without losing generality, we choose the small test set provided by (Khot et al., 2022) which contains 2024 table-statement pairs. The tables in TabFact are much smaller than those in WikiTQ. The answers are binary with nearly equal proportions.
Evaluation Metrics. We adopt the evaluator in Binder (Cheng et al., 2023) to avoid the situation where program executions are likely to be semantically correct but fail to match the gold answer exactly. For example, when the SQL outputs 1/0 for yes/no questions, the Binder evaluator will conduct a pre-matching check and translate it to boolean values first. It will avoid mismatches of correct prediction and golden answer pairs like 7.45 and 7.4499, 4 years and 4, etc. Afterwards, we adopt accuracy as our evaluation metric.
Baselines. We evaluate AutoPrep using two categories of methods: (a) TQA w/o Data Prep: Zero-Shot QA and Few-Shot QA (Chen, 2022), as well as Text-to-SQL (Rajkumar et al., 2022); (b) TQA with Data Prep: Binder (Cheng et al., 2023) and ReAcTable (Zhang et al., 2024). We further describe them below.
  • Zero-Shot and Few-Shot QA (Chen, 2022). Zero-Shot QA directly asks an LLM to generate the answer based on the serialized table and question. To improve the performance, Few-Shot QA further provides several demonstrations to enhance the In Context Learning ability of the LLM. We implement these two methods with the prompt provided by (Cheng et al., 2023).
  • Text-to-SQL (Rajkumar et al., 2022). Text-to-SQL is a natural code generation method for the TQA tasks. Consider a serialized table and a question, it first generates the SQL program and then executes it to get the final answer from the table for the question.
@@
  • Binder (Cheng et al., 2023). Binder further enhances the Text-to-SQL method by integrating LLMs into SQL programs. It uses LLMs to incorporate external knowledge bases and directly answer questions that are difficult to resolve using SQL alone. We utilize the code provided by (cod, 2023).

### v2

@@ 3. An Overview of AutoPrep
    Multiple Programmer Agents (i.e., [Uncaptioned image] Augmenter): the Programming stage. It directs the LLM to generate low-level implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$. Besides, it is also tasked for code debugging if any execution errors occur.
@@
    An [Uncaptioned image] Executor Agent: the Executing stage. It executes the given low-level code and reports error messages if any code bugs occur.
@@
  • $T^{\prime}=$Filter($T$, [Date, Country, Medal]) that filters out column Age, which is irrelevant to the question;
@@
(c) The Executing stage: an Executor agent iteratively refines the generated code if any error occurs.
@@
The Planner Agent. The key challenge here is how to suggest question-aware operations that address specific NL questions. Specifically, even for the same table, different NL questions may require not only different data prep operations but also varying sequences of those operations. For example, the second Normalize operation might not be necessary if “in February” is not part of the question. To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent. This method translates the natural language question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch. Compared to existing CoT methods (Wei et al., 2022), which break down questions into sub-questions, our approach better captures the relationships between questions and tables. More details of the CoC method are given in Section 4.
The Programmer Agents. The key challenge is that a given high-level operation can have multiple executable code alternatives (e.g., Python functions), and the difference in outcomes between the best and worst options can be substantial. For example, the Augmenter agent may generate an overly generic regular expression that extracts countries based on parentheses. Unfortunately, this code fails to correctly process “Dav. ITA”, which is formatted differently from other values. To tackle this challenge, we develop a tool-augmented approach that enhances the LLM’s code generation capabilities by utilizing predefined API functions. We establish a search space of API functions for all high-level operations, create an algorithm to identify the most suitable function for a given operation, and optimize it using a trial-and-error mechanism. More details of our tool-augmented approach are discussed in Section 5.
@@
4. The Planner
Given a question $Q$ posed over a table $T$, Planner aims to generate a logical plan that outlines a sequence of high-level operations needed to prepare the table for answering the question.
@@ 4.1. A Direct Prompting Method
The most common way to generate a logical plan is to directly prompt an LLM using a typical in-context learning approach. The inputs are a question $Q$, a table $T$, a set $\Sigma$ of specifications for each operation type, and an LLM $\theta$. Here, each specification $\sigma\in\Sigma$ describes the purpose of an operation type, e.g., “an Augment operation augments a new column for a table based on existing columns, in response to the specific needs of a question”. The output of the algorithm is a set $O$ of high-level operations like those shown in Figure 4, e.g., Normalize(“Cast to INT”, Medal).
@@ 4.2. A Chain-of-Clauses Method
To address this challenge, we propose a Chain-of-Clauses (CoC) reasoning method that decomposes the entire process of logical plan generation into two phases, as shown in Figure 5(b).
@@
Compared with existing CoT methods (Wei et al., 2022), which simply break down questions into sub-questions, our approach is more effective for logical plan generation. First, our method decomposes the question into a set of analysis steps over table, formalized as an SQL-like Analysis Sketch, which simplifies the task of logical plan generation. More importantly, our method jointly considers each analysis step along with the corresponding relevant columns (instead of the whole table) to prompt the LLM, effectively capturing the relationships between the question and the data.
@@
where ${\tt agg}(A)$ is an aggregation function (e.g., SUM and AVG) over column $A$, ${\tt Pred}(A_{i})$ denotes a predicate (i.e., filtering condition) over column $A_{i}$, such as Date LIKE02-%. Note that $f(\{A_{i}\})$ is a user-defined function (UDF) that maps existing columns $\{A_{i}\}$ in the table into a new column. Figure 5(b) shows an example Analysis Sketch with a UDF function that specifies a new column Country, i.e., $f({\tt Cyclist})$ AS Country. Obviously, this UDF is introduced because the Analysis Sketch contains a GROUP BY Country clause.
@@
  • $f({\tt Cyclist})$ AS Country: this clause and relevant columns are used to prompt the LLM to generate an Augment operation.
@@
  • Date LIKE02-%’: this clause and the values in column Date are used to prompt the LLM to generate a Normalize.
Finally, the algorithm generates a Filter operation that only selects columns relevant to the Analysis Sketch.
5. The Programmer & Executor
Programmer agents translate a high-level logical plan into a physical plan by generating low-level code, which is then passed to an Executor agent for code execution and interactive debugging. A straightforward approach is to prompt an LLM with the logical plan using in-context learning with several exemplars, asking the LLM to generate code for each high-level operation in the plan. The code is then executed iteratively, and if any runtime errors occur, the Programmer is prompted with the error messages for debugging.
However, the above method may have limitations, as the generated code may be overly generic and unable to effectively address the heterogeneity challenge of the tables. Specifically, many tables originate from web sources or real-world scenarios, where values have diverse syntactic (e.g., “19-Oct” and “9/14” in $T_{6}$) or semantic formats (e.g., “ITA” and “Italia” in $T_{4}$), making it difficult to generate code customized to these tables. For example, consider table $T_{4}$ in Figure 2(b). Given a Normalize operation to standardize country formats, a generic function from an existing Python library may not be sufficient to transform country names (e.g., “Italia”) to their ISO codes (e.g., “ITA”). In such cases, customized functions, such as the clean_country function from an external library dataprep (Peng et al., 2021), are needed to address specific requirements of code generation.
To address these limitations, we introduce a tool-augmented method for the Programmer agents, enhancing the LLM’s code generation capabilities by utilizing pre-defined API functions, referred to as the function pool in this paper.
5.1. Tool-Augmented Method: Key Idea
@@
Figure 6 provides an overview of our tool-augmented method for physical plan generation and execution. Given a table $T$ and a set $O=\{o_{1},\ldots,o_{|O|}\}$ of high-level operations, the method iteratively generates and executes low-level code for each individual operation $o_{i}\in O$, thus generating a sequence of intermediate tables $T_{1}$ (i.e., the input $T$), $T_{2}$, …$T_{|O|+1}$.
@@
Code Execution and Debugging. Given the function generated in the previous step, the Executor agent then applies the function over table $T_{i}$. If any bugs occur during execution, it captures and summarizes error messages and returns to corresponding Programmer agent. For instance, as shown in Figure 6, the first lambda function for the argument func only extracts countries based on parentheses, which may produce incorrect results for the value “Dav.ITA”, as it is formatted differently from other values. In this case, based on the execution results, the Executor agent records the error log and examines relevant data to summarize reasons, which will be passed to the corresponding Programmer to modify the code. After the execution and debugging process, the method produces a new intermediate table $T_{i+1}$ and proceeds to the next operation, $o_{i+1}$, i.e., a Normalize operation in Figure 6.
@@ 5.2. Search Space of Function Pools
This section presents a search space of function pools for different types of high-level operations formalized in the current AutoPrep.
@@
  • $\mathtt{extract}$. This function extracts a substring from a column to generate a new column. It has two arguments: ${\tt column}$, representing the name of the source column, and ${\tt func}$, representing the substring extraction function. Using table $T_{1}$ from Figure 2(a) as an example, to extract the country code, we would set ${\tt column}$ to Cyclist and ${\tt func}$ to a lambda function lambda x: re.search(r’(.?)(.*?)( . ∗ ? )’,x).group(1).
  • $\mathtt{calculate}$. The function generates a new column through arithmetic operations of existing columns. It has two arguments: ${\tt columns}$, representing a list of source columns and ${\tt func}$, representing a lambda function with input to be a dictionary of values, since we may perform calculations on multiple columns. For example, to generate a new column GrowthRate for $T_{2}$ in Figure 2(a), we set ${\tt columns}$ to [’2012’, ’2013’] and ${\tt func}$ to a lambda function lambda x: (x[’2013’]-x[’2012’])/x[’2012’].
  • $\mathtt{map\mbox{-}to\mbox{-}boolean}$ and $\mathtt{concatenate}$. Both functions generate a new column from multiple columns. The difference is that one generates boolean values, while the other concatenates strings. Both operators take two arguments, ${\tt columns}$ and ${\tt func}$, similar to those in the $\mathtt{calculate}$ operator.
@@
  • $\mathtt{to\mbox{-}numerical}$. This function standardizes a column to a numerical type, such as int or float. To use it, two arguments need to be specified: ${\tt column}$, representing the source column, and ${\tt func}$, representing a lambda function. For example, given the table $T_{5}$ in Figure 2(b), we use this function to standardize the Females column to integers, with ${\tt column}$ as Females and ${\tt func}$ as lambda x: int(x.replace(’,’, ’’)).
@@
  • $\mathtt{clean\mbox{-}string}$. To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$. It has two arguments: ${\tt column}$ and ${\tt trans\_dict}$. The $\tt trans\_dict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$. Given the original table $T_{4}$ in Figure  2(b), we set $\tt trans\_dict$ to {’Italia’: ’ITA’}.
@@
Discussions. Our search space is extensible, allowing additional functions or external APIs to be easily incorporated into the function pools, such as specific data prep functions for various semantic types developed in the dataprep library (Peng et al., 2021).
@@ 6.1. Experimental Setup
Datasets. We evaluate our proposed framework AutoPrep using two well-adopted benchmarking datasets WikiTQ (Pasupat and Liang, 2015) and TabFact (Chen et al., 2019) and a very recent benchmark TableBench (Wu et al., 2024). The statistics of the datasets are shown in Table 1.
(1) WikiTQ contains complex questions annotated by crowd workers based on diverse Wikipedia tables. WikiTQ comprises 17,689 question-answer pairs in the training set and 4,344 in the test set with significant variation in table size. The answers in WikiTQ can be categorized into three forms: (i) string from a cell in the table, (ii) string from analyzing, and (iii) list of cells in the table.
(2) TabFact provides a collection of Wikipedia tables, along with manually annotated NL statements. One requires to deduce the relations between statements and tables as “true” (a statement is entailed by a table) or “false” (a statement if refuted by a table). To reduce the experimental cost without losing generality, we choose the small test set provided by (Ye et al., 2023) which contains 2024 table-statement pairs. The tables in TabFact are much smaller than those in WikiTQ. The answers are binary with nearly equal proportions.
(3) TabBench provides an evaluation of table question answering capabilities within four major categories, including multi-hop fact checking (FC), multi-hop Numerical Reasoning (NR), Trend Forecasting and Chart Generation. We evaluate AutoPrep on TabBench-FC and TabBench-NR, where one requires to conduct more complex and multi-hop reasoning over tables for answering a question. The answers in TabBench could be a string or list of items.
Baselines. We consider the following two categories of baselines:
(1) TQA Methods w/o Data Prep directly generate the answer or utilize programs to extract the answer from the table without considering data prep. We implement four primary TQA baselines.
End2EndQA (End2End) (Chen, 2022) utilizes the in-context learning abilities of LLMs to generate the answer for TQA task based on the supervision of human-designed demonstrations. We implement End2End method with prompt and demonstrations provided by (Cheng et al., 2023).
Chain-of-Thought (CoT) (Wei et al., 2022) prompts LLMs to generate the reasoning process step-by-step before generating the final answer. We implement CoT with the prompt provided by (Chen, 2022).
NL2SQL (Rajkumar et al., 2022) first translates the question into a SQL program and then executes it to get the final answer from the table for the question. We use the prompt from (Cheng et al., 2023) to implement NL2SQL.
NL2Py uses Python code to process and reason over the tables. To construct the prompt for NL2Py, we use TQA instances in NL2SQL prompt and manually write the Python code to process the table and generate the final answer.
@@
TableBench $886$ 2$\sim$212 2$\sim$20 string / list (31.49%)
(2) TQA Methods with Data Prep. We also investigate four SOTA TQA methods that consider data prep tasks in their question answering process.
ICL-Prep. uses few-shot in-context learning (ICL) demonstrations to guide LLMs in generating programs for data prep, as shown in Figure 3. Subsequently, we employ NL2SQL to extract answers from the cleaned tables.
Dater (Ye et al., 2023) addresses the TQA task by decomposing the table and question. It first selects relevant columns and rows to obtain a sub-table and then decomposes the origin question into sub-questions. Dater answers these sub-questions based on the sub-tables to generate the final answer. We use code in (cod, 2023b) for implementation.
@@
Binder (Cheng et al., 2023) enhances the NL2SQL method by integrating LLMs into SQL programs. It uses LLMs to incorporate external knowledge bases and directly answer questions that are difficult to resolve using SQL alone. We utilize the original code provided by (cod, 2023a).